OpenStudy (anonymous):

PLEASE HELP. i need a trick & answer!? GIVING MEDAL Solve m + 11 – w = n for m.

OpenStudy (anonymous):

m=18

OpenStudy (anonymous):

M could be anything because we don't know W

OpenStudy (anonymous):

a. m = –11 + w + n c. m = + 11 + w + n b. m = –11 – w – n d. m = –11 + w – n

OpenStudy (ipwnbunnies):

Bring the 11 and the w to the right side of the equation to get 'm' by itself.